# DINOv2: Learning Robust Visual Features without Supervision

### v1

@@ 6.2 Pretraining Data Source
Training Data INet-1k Im-A ADE-20k Oxford-M
INet-22k 85.9 73.5 46.6 62.5
INet-22k $\setminus$ INet-1k 85.3 70.3 46.2 58.7
Uncurated data 83.3 59.4 48.5 54.3
LVD-142M 85.8 73.9 47.7 64.6
Table 2: Ablation of the source of pretraining data. We compare the INet-22k dataset that was used in iBOT to our dataset, LVD-142M. Each model is trained for the same number of iterations, that is smaller than in our final run. Pretraining on LVD-142M maintains the performance over INet-1k while leading to models that perform better in other domains.
The most salient observation is that training on a curated set of images works better on most benchmarks than training on uncurated data. This confirms the benefit of curating data, even in the case of self-supervised pretraining. When compared with models trained on ImageNet-22k, training on LVD-142M is also superior on all the benchmarks but ImageNet-1k. This confirms that training on a more diverse set of images improves the quality of the features in domains that are not covered by this dataset. Overall, the conclusion of this ablation is that our dataset provides a good balance of different types of images that leads to the best performance overall.
@@ 6.5 Impact of Knowledge Distillation
For small architectures, we distill larger models instead of training them from scratch. We use the distillation procedure described in Sec. 5. We evaluate the effectiveness of this approach by comparing a ViT-L/14 trained from scratch with one distilled from a ViT-g/14 over 12 benchmarks in Fig. 5. We also report the performance of the ViT-g/14 used for distillation as a topline. The distilled model outperforms the one trained from scratch on 10 out of 12 benchmarks, validating our pretraining approach for small models.
@@ 6.6 Impact of Resolution
We measure the impact of changing the resolution during the pretraining on the performance of image and patch-level features. We consider models trained from scratch using a fixed resolution of either $224\times 224$ or $416\times 416$, and a model trained from scratch at $224\times 224$, then resumed for 10k more iterations at $416\times 416$. High-resolution training is compute-intensive, so we conduct this ablation on a small setup: a ViT-L/16 trained on ImageNet1k. In Fig. 6, we report the performance of a linear probe on ImageNet-1k and ADE-20k, evaluated at various resolutions. The model trained on high-resolution images performs best across resolutions, but this comes at a high cost: training at $416$ is approximate $3\times{}$ more compute-intensive than training at $224$. On the other hand, training at high resolution for only 10k iterations at the end of the training is almost as good and only requiring a fraction of the compute. As a consequence, we include this step at the end of the training rather than training at a high resolution from scratch.
@@ 7 Results
Baselines. In our comparisons, we use two kinds of models as baselines. We compare to the best performing self-supervised models that are openly available. First, we run our evaluations for MAE (He et al., 2021), DINO (Caron et al., 2021), SEERv2 (Goyal et al., 2022a), MSN (Assran et al., 2022), EsViT (Li et al., 2022a), Mugs (Zhou et al., 2022) and iBOT (Zhou et al., 2021). When several architectural variants were proposed for a given method, we report results for the one that leads to best top-1 accuracy on ImageNet-1k. Second, we report performance of open-source weakly-supervised models such as CLIP (Radford et al., 2021), OpenCLIP (Ilharco et al., 2021), and SWAG (Singh et al., 2022). When evaluating models on ImageNet-1k, we report the performance for each of the aforementioned methods. For all other evaluations, we report the four best-performing models amongst SSL ones. Also, for reference, we report the best performing OpenCLIP-G for weakly-supervised ones.
@@ 7.1 ImageNet Classification
OpenCLIP ViT-H/14 LAION 81.7 84.4 88.4 75.5
OpenCLIP ViT-G/14 LAION 83.2 86.2 89.4 77.2
@@
We question if the ability of our models to produce high quality frozen features impact their performance when finetuned with supervision on a specific dataset. While this is not core to this paper, this experiment is indicative of whether we have involuntarily specialized our models to the setting of linear evaluations of frozen features. To run this sanity check, we apply the finetuning pipeline from Touvron et al. (2022), without tweaking hyper-parameters. In Table 5, we show that the Top-1 accuracy on the validation set of ImageNet-1k improves by more than $2\%$ when the backbone is finetuned. This is true both when using models at resolution $224$ and $448$. Further gains can be obtained by tuning the hyper-parameters of the finetuning, but this is beyond the goal of this sanity check. Nonetheless, our best finetuned performance ($88.9\%$) is only a couple of percent below ($2.2\%$) the absolute state of the arts ($91.1\%$), obtained by Chen et al. (2023). As DINOv2 leads to features that are strong in both the linear and finetuning settings, a strong property of our approach is that finetuning is optional.
@@
To complement our study, and probe the generalization of our features, we evaluate our ImageNet-1k models trained with linear classification heads on domain generalization benchmarks. We use the best performing linear classifier as described above and simply run inference on those benchmarks. Please note that most results in the litterature are obtained with models that are finetuned end-to-end on ImageNet-1k. We show the result of this experiment in Table 6. When comparing with state-of-the-art SSL methods, our models shows drastically better robustness ($29.6\%$ on A, $22.1\%$ on R and $23.0\%$ on Sketch compared to iBOT). Our model also improves upon the best weakly-supervised model on ImageNet-A while lagging behind on R and Sketch.
@@ 7.4 Dense Recognition Tasks
In a final experiment, we freeze our backbone, and plug it into a ViT-Adapter Chen et al. (2022) with a Mask2former head (Cheng et al., 2022). We tune the weights of the adapter and head, but keep the backbone frozen: only a fraction of the weights are tuned, keeping the training procedure lightweight. We reach $60.2$ mIoU on ADE20k, close to the competitive state of the art, standing at 62.9 mIoU (Wang et al., 2022).
@@
Table 10: Semantic segmentation on ADE20K, CityScapes and Pascal VOC with frozen features and a linear classifier (lin.) and with multiscale (+ms). The absolute state of the art – from Wang et al. (2022), Liu et al. (2021) and Chen et al. (2018) respectively – are mentioned at the top of the Table. For reference, using the Mask2Former pipeline (Steiner et al., 2021) with a ViT-Adapter (Chen et al., 2022) on top of our frozen ViT-g/14 backbone gives 60.2 mIoU on ADE-20k.
@@ 8.1 Geographical Fairness
We evaluate geographical fairness on the Dollar Street dataset introduced in De Vries et al. (2019) using the evaluation protocol of Goyal et al. (2022b). This benchmark compares performance across countries and income levels. It contains 16,073 images from 289 households across 54 countries. The task is to recognize 94 concepts that vary visually between households based on income or location. In Table 12, we compare our model with SEERv2 (Goyal et al., 2022a), a model trained on a geographically diverse set of images. Our model is slightly fairer across regions and incomes than the SEERv2 model and significantly better than the supervised baseline reported by Goyal et al. (2022a). However, we still observe a significant difference between regions, particularly in Africa, where our model performance drops by 25.7% compared to Europe. This show that our model is still biased toward Western countries. Similarly, our model performs significantly better on high-income households than low-income ones, with a difference of 31.7%. Despite improvements, we observe significant biases in our models toward wealthy households from Western countries.
@@ 9 Estimating the Environmental Impact of Training our Models
Additionally, we estimate the footprint of the whole project to be between 0.50.5k and 11k tCO2eq using the same grid as presented above. This carbon footprint represents in the order of 200200k GPU-days. The primary sources of emissions are the self-supervised pre-trainings of the models. For example, a single pre-training of a ViT-g model (22k GPU-hours) emits 3.7 tons of CO2eq, while a finetuning on ImageNet-1k (1k GPU-hours) emits 0.2 tons. This estimate only considers the GPUs’ electricity consumption and ignores other emissions, such as their manufacturing and disposal.
@@ 10 Future work and Discussion
In this work, we present DINOv2, a new series of image encoders pretrained on large curated data with no supervision. This is the first SSL work on image data that leads to visual features that close the performance gap with (weakly) supervised alternatives across a wide range of benchmarks and without the need for finetuning. A few properties emerge from these models, such as an understanding of object parts and scene geometry regardless of the image domains. We expect that more of these properties will emerge at larger scales of models and data, akin to instruction emergence in large language models, and plan to continue scaling along these axes. This paper also demonstrates that these visual features are compatible with classifiers as simple as linear layers - meaning the underlying information is readily available. In future work, we plan to leverage this ability to train a a language-enabled AI system that can process visual features as if they were word tokens, and extract the required information to ground the system.
@@ Acknowledgments.
We thank Mathilde Caron for initial discussions that led to this work. We thank Olivia Joulin for the horse drawing used in Fig. 10. We also thank the rest of FAIR and Meta AI for feedback on this work through the entire project.
@@ A.2 Image similarity
$m(s,r)=\text{cosine-similarity}\left(f\left(s\right),f\left(r\right)\right)=\frac{f(s).f(r)}{\lVert f(s)\rVert_{2}\lVert f(r)\rVert_{2}}$
@@ A.4 Retrieval
Table 15: Composition of our LVD-142M dataset. We report the list of datasets and associated splits used to build the dataset, how they were included (as is without retrieval or via sample-based or cluster-based retrieval). For retrievals, we indicate the actual number of retrieved images and the final number included in the dataset.
@@ B.3 Linear probing evaluation
Appendix C List of benchmarks used for evaluations
We show in Table 18 the list of benchmarks and datasets used for evaluation.
Dataset name Task Citation
ImageNet-1k Image Classification (Russakovsky et al., 2015)
ImageNet-V2 Image Classification (Recht et al., 2019)
ImageNet-ReaL Image Classification (Beyer et al., 2020)
ImageNet-A Image Classification (Djolonga et al., 2021)
ImageNet-C Image Classification (Hendrycks & Dietterich, 2019)
ImageNet-Rendition Image Classification (Hendrycks et al., 2021)
ImageNet-Sketch Image Classification (Wang et al., 2019)
Food-101 Image Classification (Bossard et al., 2014)
CIFAR-10 Image Classification (Krizhevsky et al., 2009)
CIFAR-100 Image Classification (Krizhevsky et al., 2009)
SUN397 Image Classification (Xiao et al., 2010)
StanfordCars Image Classification (Krause et al., 2013)
FGVC-Aircraft Image Classification (Maji et al., 2013)
Pascal VOC 2007 Image Classification (Everingham et al., 2015)
Describable Textures Image Classification (Cimpoi et al., 2014)
Oxford Pets Image Classification (Parkhi et al., 2012)
Caltech101 Image Classification (Fei-Fei et al., 2004)
Oxford Flowers Image Classification (Nilsback & Zisserman, 2008)
CUB200 Image Classification (Welinder et al., 2010)
iNaturalist 2018 Image Classification (Van Horn et al., 2018)
iNaturalist 2021 Image Classification (Van Horn et al., 2021)
Places-205 Image Classification (Zhou et al., 2014)
UCF101 Video Classification (Soomro et al., 2012)
Kinetics-400 Video Classification (Kay et al., 2017)
Something-Something-V2 Video Classification (Goyal et al., 2017)
Revisiting-Paris Image Retrieval (Radenović et al., 2018a)
Revisiting-Oxford Image Retrieval (Radenović et al., 2018a)
Met Image Retrieval (Ypsilantis et al., 2021)
Amstertime Image Retrieval (Yildiz et al., 2022)
ADE20k Image Segmentation (Zhou et al., 2017)
Cityscapes Image Segmentation (Cordts et al., 2016)
Pascal VOC 2012 Image Segmentation (Everingham et al., 2015)
NYU-Depth V2 Monocular Depth Estimation (Silberman et al., 2012)
KITTI Monocular Depth Estimation (Geiger et al., 2013)
SUN-RGBD Monocular Depth Estimation (Song et al., 2015)
DollarStreet Fairness Analysis (De Vries et al., 2019)
Casual Conversations Fairness Analysis (Hazirbas et al., 2021)
Table 18: List of datasets used for evaluation.

### v2

@@ 6.2 Pretraining Data Source
Training Data INet-1k Im-A ADE-20k Oxford-M iNat2018 iNat2021 Places205
INet-22k 85.9 73.5 46.6 62.5 81.1 85.6 67.0
INet-22k $\setminus$ INet-1k 85.3 70.3 46.2 58.7 80.1 85.1 66.5
Uncurated data 83.3 59.4 48.5 54.3 68.0 76.4 67.2
LVD-142M 85.8 73.9 47.7 64.6 82.3 86.4 67.6
Table 2: Ablation of the source of pretraining data. We compare the INet-22k dataset that was used in iBOT to our dataset, LVD-142M. Each model is trained for the same number of iterations, that is smaller than in our final run, without high-resolution adaptation. Pretraining on LVD-142M maintains the performance over INet-1k while leading to models that perform better in other domains.
The most salient observation is that training on a curated set of images works better on most benchmarks than training on uncurated data. This confirms the benefit of curating data, even in the case of self-supervised pretraining. When compared with models trained on ImageNet-22k, training on LVD-142M is also superior on all the benchmarks but ImageNet-1k. This confirms that training on a more diverse set of images improves the quality of the features in domains that are not covered by ImageNet-22k. We also see that training on our curated data increases the performances on domains that are not used for the curation process (INaturalist 2018, 2021 and Places205), proving that scale and diversity can benefit unseen domains. Overall, the conclusion of this ablation is that our dataset provides a good balance of different types of images that leads to the best performance overall.
@@ 6.5 Impact of Knowledge Distillation
For small architectures, we distill larger models instead of training them from scratch. We use the distillation procedure described in Sec. 5. We evaluate the effectiveness of this approach by comparing a ViT-L/14 trained from scratch with one distilled from a ViT-g/14 over 12 benchmarks in Fig. 5. We also report the performance of the ViT-g/14 used for distillation as a topline. The distilled model outperforms the one trained from scratch on all 12 benchmarks, validating our pretraining approach for small models.
@@ 6.6 Impact of Resolution
We measure the impact of changing the resolution during the pretraining on the performance of image and patch-level features. We consider models trained from scratch using a fixed resolution of either $224\times 224$ or $416\times 416$, and a model trained from scratch at $224\times 224$, then resumed for 10k more iterations at $416\times 416$. High-resolution training is compute-intensive, so we conduct this ablation on a small setup: a ViT-L/16 trained on ImageNet1k. In Fig. 6, we report the performance of a linear probe on ImageNet-1k and ADE-20k, evaluated at various resolutions. The model trained on high-resolution images performs best across resolutions, but this comes at a high cost: training at $416$ is approximately $3\times{}$ more compute-intensive than training at $224$. On the other hand, training at high resolution for only 10k iterations at the end of the training is almost as good and only requiring a fraction of the compute. As a consequence, we include this step at the end of the training rather than training at a high resolution from scratch.
@@ 7 Results
Baselines. In our comparisons, we use two kinds of models as baselines. We compare to the best performing self-supervised models that are openly available. First, we run our evaluations for MAE (He et al., 2022), DINO (Caron et al., 2021), SEERv2 (Goyal et al., 2022a), MSN (Assran et al., 2022), EsViT (Li et al., 2022a), Mugs (Zhou et al., 2022b) and iBOT (Zhou et al., 2022a). When several architectural variants were proposed for a given method, we report results for the one that leads to best top-1 accuracy on ImageNet-1k. Second, we report performance of open-source weakly-supervised models such as CLIP (Radford et al., 2021), OpenCLIP (Ilharco et al., 2021; Cherti et al., 2023), and SWAG (Singh et al., 2022). When evaluating models on ImageNet-1k, we report the performance for each of the aforementioned methods. For all other evaluations, we report the four best-performing models amongst SSL ones. Also, for reference, we report the best performing OpenCLIP-G for weakly-supervised ones.
@@ 7.1 ImageNet Classification
OpenCLIP ViT-H/14 LAION-2B 81.7 84.4 88.4 75.5
OpenCLIP ViT-G/14 LAION-2B 83.2 86.2 89.4 77.2
@@
We question if the ability of our models to produce high quality frozen features impact their performance when finetuned with supervision on a specific dataset. While this is not core to this paper, this experiment is indicative of whether we have involuntarily specialized our models to the setting of linear evaluations of frozen features. To run this sanity check, we apply the finetuning pipeline from Touvron et al. (2022), without tweaking hyper-parameters. In Table 5, we show that the Top-1 accuracy on the validation set of ImageNet-1k improves by more than $2\%$ when the backbone is finetuned. This is true both when using models at resolution $224$ and $448$. Further gains can be obtained by tuning the hyper-parameters of the finetuning, but this is beyond the goal of this sanity check. Nonetheless, our best finetuned performance ($88.9\%$) is only a couple of percent below ($2.2\%$) the absolute state of the arts ($91.1\%$), obtained by Chen et al. (2023a). As DINOv2 leads to features that are strong in both the linear and finetuning settings, a strong property of our approach is that finetuning is optional.
@@
To complement our study, and probe the generalization of our features, we evaluate our ImageNet-1k models trained with linear classification heads on domain generalization benchmarks. We use the best performing linear classifier as described above and simply run inference on those benchmarks. Please note that most results in the literature are obtained with models that are finetuned end-to-end on ImageNet-1k. We show the result of this experiment in Table 6. When comparing with state-of-the-art SSL methods, our models shows drastically better robustness ($29.6\%$ on A (Hendrycks et al., 2021b), $22.1\%$ on R (Hendrycks et al., 2021a) and $23.0\%$ on Sketch (Wang et al., 2019) compared to iBOT). Our model also improves upon the best weakly-supervised model on ImageNet-A while lagging behind on R and Sketch.
@@ 7.4 Dense Recognition Tasks
Frozen backbone in a SOTA pipeline.
In a final experiment, we freeze our backbone, and plug it into a ViT-Adapter Chen et al. (2023b) with a Mask2former head (Cheng et al., 2022). We tune the weights of the adapter and head, but keep the backbone frozen, meaning 66% of the weights are frozen. This allows for a lighter segmentation training than full end-to-end fine-tuning. With this setup, we reach $60.2$ mIoU on ADE20k, close to the competitive state of the art, standing at 62.9 mIoU (Wang et al., 2022). Although our setup for this experiment doesn’t makes use of the optimisations described in Sec. 5, the segmentation training in this experiment took 28 hours on 16 V100 GPUs.
@@
Table 10: Semantic segmentation on ADE20K, CityScapes and Pascal VOC with frozen features and a linear classifier (lin.) and with multiscale (+ms). The absolute state of the art – from Wang et al. (2022), Liu et al. (2021) and Chen et al. (2018) respectively – are mentioned at the top of the Table. For reference, using the Mask2Former pipeline (Steiner et al., 2021) with a ViT-Adapter (Chen et al., 2023b) on top of our frozen ViT-g/14 backbone gives 60.2 mIoU on ADE-20k.
@@ 8.1 Geographical Fairness
We evaluate geographical fairness on the Dollar Street dataset introduced in De Vries et al. (2019) using the evaluation protocol of Goyal et al. (2022b). This benchmark compares performance across countries and income levels. It contains 16,073 images from 289 households across 54 countries. The task is to recognize 94 concepts that vary visually between households based on income or location. In Table 12, we compare our model with SEERv2 (Goyal et al., 2022a), a model trained on a geographically diverse set of images. Our model is slightly fairer across regions and incomes than the SEERv2 model and significantly better than the supervised baseline reported by Goyal et al. (2022a). However, we still observe a significant difference between regions, particularly in Africa, where our model performance drops by 25.7% compared to Europe. This shows that our model is still biased toward Western countries. Similarly, our model performs significantly better on high-income households than low-income ones, with a difference of 31.7%. Despite improvements, we observe significant biases in our models toward wealthy households from Western countries.
@@ 9 Estimating the Environmental Impact of Training our Models
Additionally, we estimate the footprint of the whole project to be between 0.50.5k and 11k tCO2eq using the same grid as presented above  [^3] For context, a full Boeing 777 return flight between London and New York corresponds to approximately 560 tCO2eq.. This carbon footprint represents in the order of 200200k GPU-days. The primary sources of emissions are the self-supervised pre-trainings of the models. For example, a single pre-training of a ViT-g model (22k GPU-hours) emits 3.7 tons of CO2eq, while a finetuning on ImageNet-1k (1k GPU-hours) emits 0.2 tons. This estimate only considers the GPUs’ electricity consumption and ignores other emissions, such as their manufacturing and disposal.
@@ 10 Future work and Discussion
In this work, we present DINOv2, a new series of image encoders pretrained on large curated data with no supervision. This is the first SSL work on image data that leads to visual features that close the performance gap with (weakly) supervised alternatives across a wide range of benchmarks and without the need for finetuning. We can attribute the strong performance of the DINOv2 family of models to several factors: i) an improved training recipe with better hyperparameters and regularization (Table 1), ii) a larger model scale with improved results regardless of the data used for training (Fig. 4), iii) a larger dataset (Fig. 4) and iv) the distillation process that makes smaller models benefit from the performance of the strongest ViT-g model (Fig. 5). A few properties emerge from these models, such as an understanding of object parts and scene geometry regardless of the image domains. We expect that more of these properties will emerge at larger scales of models and data, akin to instruction emergence in large language models, and plan to continue scaling along these axes. This paper also demonstrates that these visual features are compatible with classifiers as simple as linear layers - meaning the underlying information is readily available. In future work, we plan to leverage this ability to train a a language-enabled AI system that can process visual features as if they were word tokens, and extract the required information to ground the system.
@@ Acknowledgments.
We thank Mathilde Caron for initial discussions that led to this work. Julien Mairal was supported by the ERC grant number 101087696 (APHELAIA project) and by ANR 3IA MIAI@Grenoble Alpes (ANR-19-P3IA-0003). We thank Olivia Joulin for the horse drawing used in Fig. 10. We thank Madeleine and Léon for posing for Fig. 8 We also thank the rest of FAIR and Meta AI for feedback on this work through the entire project.
@@ A.2 Image similarity
$m(s,r)=\text{cosine-similarity}\left(f\left(s\right),f\left(r\right)\right)=\frac{f(s)\cdot{}f(r)}{\lVert f(s)\rVert_{2}\lVert f(r)\rVert_{2}}$
@@ A.4 Retrieval
Table 15: Composition of our LVD-142M dataset. We report the list of datasets and associated splits used to build the dataset, how they were included (as is without retrieval or via sample-based or cluster-based retrieval). For retrievals, we indicate the actual number of retrieved images and the final number included in the dataset. We chose to include as many datasets as possible in the pretraining data in order to cover as many domains as possible. We kept a few datasets aside in order to evaluate performance outside of the pretraining domain. More details about dataset usages can be found in Table 18.
@@ B.3 Linear probing evaluation
Appendix C List of Datasets used
We show in Table 18 the list of benchmarks and datasets used and their purposes.
Dataset Pretraining (as is) Retrieving pretraining data Eval. Task Citation
ImageNet-1k Classif.  (Russakovsky et al., 2015)
ImageNet-22k  (Deng et al., 2009)
ImageNet-V2 Classif.  (Recht et al., 2019)
ImageNet-ReaL Classif.  (Beyer et al., 2020)
ImageNet-A Classif.  (Hendrycks et al., 2021b)
ImageNet-C Classif.  (Hendrycks & Dietterich, 2019)
ImageNet-R Classif.  (Hendrycks et al., 2021a)
ImageNet-Sk. Classif.  (Wang et al., 2019)
Food-101 Classif.  (Bossard et al., 2014)
CIFAR-10 Classif.  (Krizhevsky et al., 2009)
CIFAR-100 Classif.  (Krizhevsky et al., 2009)
SUN397 Classif.  (Xiao et al., 2010)
StanfordCars Classif.  (Krause et al., 2013)
FGVC-Aircraft Classif.  (Maji et al., 2013)
VOC 2007 Classif.  (Everingham et al., 2010)
DTD Classif.  (Cimpoi et al., 2014)
Oxford Pets Classif.  (Parkhi et al., 2012)
Caltech101 Classif.  (Fei-Fei et al., 2004)
Flowers Classif.  (Nilsback & Zisserman, 2008)
CUB200 Classif.  (Welinder et al., 2010)
iNaturalist 2018 Classif.  (Van Horn et al., 2018)
iNaturalist 2021 Classif.  (Van Horn et al., 2021)
Places-205 Classif.  (Zhou et al., 2014)
UCF101 Video  (Soomro et al., 2012)
Kinetics-400 Video  (Kay et al., 2017)
SSv2 Video  (Goyal et al., 2017)
GLD v2  (Weyand et al., 2020)
R-Paris Retrieval  (Radenović et al., 2018a)
R-Oxford Retrieval  (Radenović et al., 2018a)
Met Retrieval  (Ypsilantis et al., 2021)
Amstertime Retrieval  (Yildiz et al., 2022)
ADE20k Seg.  (Zhou et al., 2017)
Cityscapes Seg.  (Cordts et al., 2016)
VOC 2012 Seg.  (Everingham et al., 2010)
Mapillary SLS  (Warburg et al., 2020)
NYU-Depth V2 Depth  (Silberman et al., 2012)
KITTI Depth  (Geiger et al., 2013)
SUN-RGBD Depth  (Song et al., 2015)
DollarStreet Fairness  (De Vries et al., 2019)
Casual Conv. Fairness  (Hazirbas et al., 2021)
Table 18: List of datasets used.